OpenStudy (anonymous):

cotx(cosx -1) wrting this in terms of sinx cosx/sinx * cosx-1/1/secx is what i got so far but not sure if I am in ball park or if I am off with the secx?

OpenStudy (anonymous):

im thinking my answer would be secx and that the cosx would cancel out but not sure

OpenStudy (anonymous):

Your question says to write it in terms of sin x. sec x is not in terms of sin x.....

OpenStudy (anonymous):

\[\cot x(\cos x-1)=\frac{\cos x}{\sin x}(\cos x-1)\] \[\large =\frac{\cos^2 x}{\sin x}-\frac{\cos x}{\sin x}\] \[\large =\frac{\cos^2 x-\cos x}{\sin x}\] Using the identity: \[\cos^2 x+\sin^2 x=1\] \[\cos^2 x=1-\sin^2 x\] \[\cos x =\pm \sqrt{1-\sin^2 x}\] \[\large =\frac{1-\sin^2 x-\cos x}{\sin x}\] \[\large =\frac{1-\sin^2 x\pm \sqrt{1-\sin^2 x}}{\sin x}\]

OpenStudy (anonymous):

so we factored the expressions in order to solve?

OpenStudy (anonymous):

Where does your question say solve unless this is one part of a question.

OpenStudy (anonymous):

it doesnt

OpenStudy (anonymous):

So Voila! We're done here if you understand completely what I did there. If not, then you can ask away at anything that's mind-boggling about the question.

OpenStudy (anonymous):

cos2xsinx−cosxsinx I gues is what i don't understand is how you got this

OpenStudy (anonymous):

I distributed/expanded to get rid of the brackets.

OpenStudy (anonymous):

paranthesis*

OpenStudy (anonymous):

oh ok I see it now \thank you for your assistance

OpenStudy (anonymous):

No worries.